OpenStudy (anonymous):

in a orchard there are 150 oranges trees and 250 apples trees. In simplest fraction form what is the ratio of oranges trees to apple trees

OpenStudy (inkyvoyd):

150/250=15/25=3/5

OpenStudy (inkyvoyd):

\(\Huge \frac{150}{250}=\frac{15}{25}=\frac{3}{5}\)

OpenStudy (anonymous):

Ratio of oragnge tree to apple tree = Orange tree/Apple Tree Ratio = 150/250 Both numbers are divisible by 50 so Ratio = (150/50)/(250/50) = 3/5

OpenStudy (inkyvoyd):

you'll want to express that as 3:5 or 3 to 5